# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

### v4

@@ 3 The HUCA Framework
However, early departures of EVs occur occasionally, making the departure time uncertain and complicating the selection of the best low-level actions. To tackle this challenge, inspired by the Upper Confidence Bound (UCB) algorithm [1], which adjusts action-value estimations based on uncertainty (or confidence) and thereby affects the action decision, we propose a novel augmented action-value function, $\mathcal{Q}^{i}_{l}$, for the actor network of each low-level agent $c_{i}$. The idea is to discourage higher power level discharging actions when the given expected departure time is approaching and the expected SoC is still unmet. This augmentation incorporates uncertainty into the actor network for decision-making, while isolating direct impact of uncertainties from the critic network training, thereby maintaining the robustness of the critic network. Accordingly, $\mathcal{Q}^{i}_{l}$ is designed as follows:
(3.6) $\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})\cdot\biggl{(}1-$
@@
$\hskip 51.21504pt\rho\cdot\underbrace{|\log_{2}(\bar{a}^{i}_{l}+\epsilon)|\cdot\sqrt{\frac{\max(\Delta SoC^{i}_{t},0)}{\Delta T^{i}_{t}}}}_{\text{the uncertainty factor}}\>\>\biggr{)}\Large\small\large,$
@@
where $\bar{a}^{i}_{l}\in[0,1]$ is the action chosen by low-level agent $c_{i}$. A smaller $\bar{a}^{i}$ indicates a tendency for higher discharge power, as defined in Sec. 3.2. $\Delta SoC^{i}_{t}$ represents the difference between the expected SoC $SoC^{i}_{\text{dep}}$ and the current SoC $SoC^{i}_{t}$ of the EV connected to pile $c_{i}$. $\Delta T^{i}_{t}$ is the time difference between the current time $t$ and the planned departure time $t^{i}_{\text{dep}}$. $\rho\in\mathbb{R}$ is a fixed coefficient representing the impact of uncertainty, and $\epsilon$ is a small constant. Consequently, high power discharging action is penalized when the uncertainty factor increases.
@@
Based on Eq. 3.6, the actor network $\theta^{\mu,i}_{l}$ of low-level agent $c_{i}$ updates the low-level policy $\mu^{i}_{l}$ with gradient descent as follows:
(3.7) $\nabla_{\theta^{\mu,i}_{l}}J(\mu^{i}_{l})=$
@@
$\hskip 14.22636pt\mathop{\mathbb{E}}_{\textbf{x},\textbf{a}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\bigl{[}\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(a^{i}_{l}|s^{i}_{l})\nabla_{a^{i}_{h}}\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})|_{a^{i}_{l}=\mu^{i}_{l}(s^{i}_{l})}\bigr{]}.$
3.2.3 Continuous actions and the reward function for low-level agents.
@@
Let $a^{i}_{l}\in[0,1]$ be the action selected by a low-level agent $c_{i}$. To entangle low-level action space with the high-level discrete decision $a^{\text{disc}}_{h}$, the continuous action $\bar{a}^{i}_{l}$ is formulated as follows:
@@
(3.8) $\bar{a}^{i}_{l}=\left\{\begin{matrix}\delta+\delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=1$}\\ \delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=0$}\end{matrix}\right.,$
@@
(3.9) $\mathcal{P}^{\text{opt}}_{i,t}=\bar{a}^{i}_{l}\cdot(\mathcal{P}^{\max}_{i,t}-\mathcal{P}^{\min}_{i,t})+\mathcal{P}^{\min}_{i,t}.$
Based on the selected action $a^{i}_{t}$ and the optimal power decision $\mathcal{P}^{\text{opt}}_{i,t}$ at each time slot, the reward function $r_{l}(s^{i}_{l},a^{i}_{l})$ (abbreviated as $r^{i}_{l}$) for the low-level agent (pile) $c_{i}$ is formulated based on the charging cost at the current time slot $t$ and the difference between the current SoC and the expected SoC at the EV’s departure connected to $c_{i}$, defined as follows:
(3.10) $r_{l}=\omega\cdot(-\mathcal{P}^{\text{opt}}_{i,t}\cdot p_{t})+(-|SoC^{i}_{t}-SoC^{i}_{\text{dep}}|),$
where $p_{t}$ and $SoC^{i}_{t}$ are the energy price and the SoC of the EV at time slot $t$, respectively. $SoC^{i}_{\text{dep}}$ is the expected SoC at the EV’s departure. The parameter $\omega$ reflects the impact of the current charging cost.
@@
Based on Sec. 3.1 and Sec. 3.2, the critic and actor networks are optimized with minibatches of samples from the replay buffers, which follows the soft update rule. Let $\theta_{h}=\{\theta^{Q}_{h},\theta^{\mu}_{h}\}$ and $\theta^{i}_{l}=\{\theta^{Q,i}_{l},\theta^{\mu,i}_{l}\}$ represent the parameters of target critic and actor networks of the high-level agent and the low-level agent ($c_{i}$), respectively. The parameters updates are as follows:
@@
where $\tau\in[0,1]$ is the coefficient controlling the update rate. Through these optimizations, the decision-making processes of both high-level and low-level agents are refined without requiring future information. The details of optimization formulation are given in Appendix A.3.
@@ 4 Experimental Results
4.1 Dataset and Experimental Setup.
4.1.1 Dataset Description and Preprocessing.
@@
For our workplace charging system scenario, we use the following datasets: (i) Building electricity demands: the CU-BEMS dataset [17] provides office building electricity consumption. (ii) Pricing mechanism: the ComEd APIs [5] supply real-time pricing data. Also, the penalty mechanism is designed based on [18]. The data is aggregated into hourly intervals. The training period is from July 1, 2018, to August 31, 2018, with testing in the following month. Details are in Appendix A.1.
4.1.2 Simulation of EV Behaviors.
The charging information $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$ (Def. 2) of the EV is modeled using truncated normal distributions by following [10]. Table 1 shows the setting of random variables and $C^{i}$ is set to 60 kWh. Two departure scenarios are examined:
  1. 1.
    Certain departure scenario: The actual departure time $\widehat{t}^{i}_{dep}$ is the same as the expected time $t^{i}_{dep}$.
  2. 2.
    Uncertain departure scenario: The actual departure time is randomly sampled earlier than the expected time, $\widehat{t}^{i}_{dep}\in[1,t^{i}_{dep})$.
Table 1: Random variables for EV information.
Information Distribution Boundaries
Arrival time $\mathcal{N}(9,1^{2})$ $7\leq t^{i}_{\text{arr}}\leq 12$
Table 2: Performance comparison (HUCA performance) with different number of charging piles (CP). Except of the “OPT” method, the best and second best results for “Penalty Cost” ($\downarrow$) and “Total Cost” ($\downarrow$) are in bold and underlined, respectively. Cells in gray highlight the results that are greater than or equal to the median value for “SoC Maintenance” ($\uparrow$), “SoC Fulfillment” ($\uparrow$) and “User Satisfaction” ($\uparrow$).
Certain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
Penalty Cost (USD) 10 0.00 3682.85 2516.84 522.28 354.84 354.82
Total Cost (USD) 5805.28 9563.31 8426.36 6431.31 6220.33 6215.42
Penalty Cost (USD) 20 0.00 7818.52 5643.15 2271.32 1869.38 1643.47
Total Cost (USD) 5860.05 13929.11 11798.68 8413.61 7952.69 7728.57
Uncertain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
Penalty Cost (USD) 10 0.00 4671.74 1742.35 529.18 448.52 355.84
Total Cost (USD) 5691.96 10398.77 7479.04 6214.53 6087.48 5993.78
SoC Fulfillment (%) 53.77 54.05 46.04 40.34 27.10 40.77
SoC Maintenance (%) 42.52 -10.58 129.01 -120.61 11.87 45.40
User Satisfaction (%) 48.14 21.73 87.52 -40.13 19.48 43.08
Penalty Cost (USD) 20 0.00 5010.81 2002.12 966.51 931.01 354.90
Total Cost (USD) 5760.05 10823.07 7815.99 6721.82 6562.80 6013.52
SoC Fulfillment (%) 56.06 55.27 47.94 48.88 40.34 42.44
SoC Maintenance (%) 32.77 12.56 0.90 91.96 -1.16 124.41
User Satisfaction (%) 44.41 33.91 24.42 70.42 19.59 83.42
Table 3: Ablation study results for the uncertain departure scenario. (SoC fulfillment and maintenance scores are abbreviated as “Ful.” and “Main.”, respectively.)
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
10 HUCA 355.84 5993.78 40.77 % 45.40%
w/o C.A. 2015.62 7761.55 62.49% 124.01%
w/o H. 2707.71 8483.96 63.10% 31.91%
w/o Either 723.46 6393.26 44.53% -12.53%
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
20 HUCA 354.90 6013.52 42.44% 124.41%
w/o C.A. 9362.87 15309.23 69.73% 209.70%
w/o H. 3461.37 9261.72 58.61% 7.69%
w/o Either 618.46 6381.87 53.26% 43.02%
4.1.3 Baselines.
We compare our HUCA with following baselines. OPT: An oracle with full knowledge of future information serves the optimal solution of minimal total cost and minimal penalty cost. We include the single-agent (DDPG [13]) and multi-agent reinforcement learning models (IQL [22], VDN[21], and MADDPG [14]). These details are provided in Appendix A.1.4.
4.1.4 Evaluation Metrics.
  • Penalty Cost (USD) measures the cost of exceeding the contracted capacity [7], which multiplies the penalty rate with the peak amount of electricity (check Appendix A.1 for details).
4.2 Comparison Results.
As shown in Table 2, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total costs that do not rely on future information for charging control across both departure scenarios. In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs. Although the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, it still achieves a comparable results to all the other baselines. Note that combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
Among the multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost. However, the user satisfaction scores of MADDPG are much lower than those of HUCA across different numbers of charging piles. While VDN achieves higher SoC fulfillment score than HUCA with 20 charging piles, its maintenance scores are significantly lower. With 10 charging piles, VDN’s maintenance score even drops to a negative value (-120.61%), indicating that under MADDPG and VDN control, EVs’ SoCs are not maintained at arrival levels. Therefore, EVs primarily serve as power providers rather than being effectively charged, resulting in additional expenses for users.
4.3 Ablation Study and Analysis of the Uncertainty Impact.
We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either). The results are shown in Table 3. Compared to HUCA, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores. However, this also leads to a significant rise in both penalty and total costs. On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores. These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while effectively balancing energy supply between the office building and EVs.
4.4 Case Study.
To examine how HUCA balances energy supply between the building and EVs, Figure 4 visualizes the control decisions over a day. Two significant EV charging behaviors appear at 7:00 and 16:00 (red rectangle). At 7:00, both electricity demand and prices for the building are minimal (Figure 4(a)), therefore HUCA focuses on charging the EVs to reach their target SoC.At 16:00, even though there is a high demand for building energy and elevated prices, which typically would lead to transferring energy from the EVs to the building to reduce costs, HUCA opts to charge the EVs because of their high charging urgency (Figure 4(b)) using Eq. (3.6). In contrast, at 8:00 (green rectangle), HUCA discharges EVs to serve the high-usage building and avoid high pricing time simultaneously. These results manifest that HUCA effectively and dynamically balances energy supply between the building and EVs based on real-time information.
5 Related Work
5.1 Real-time Charging Control.
Guo et al. [8] combined day-ahead optimization with real-time control to address uncertainties in solar output and EV parking behaviors. Chandra Mouli et al. [16] introduced an energy management system with vehicle-to-grid (V2G) services to adjust charging control. Chopra et al. [3] conducted a cost-benefit analysis for EV charging infrastructure in parking lots. Zhao et al. [29] employed risk-based day-ahead scheduling for EV aggregators using information gap decision theory to minimize building costs. Zhang et al. [28] developed an EV load optimization model to reduce grid-level energy costs by considering peak and valley electricity prices. Chen et al. [2] devised strategies for battery energy storage systems to participate in peak load shifting, minimizing load fluctuations.
5.2 Reinforcement Learning for Charging Management.
Single-agent RL, multi-agent RL (MARL), and hierarchical MARL approaches were used to manage charging systems. Single-agent RL represented the control system [12, 15] under various considerations, such as ensuring EV battery requirements [26] and addressing charger shortage [9]. MARL modeled fine-grained cooperations [25] among EVs and buildings to tackle safety concerns [27] and power transfer overload [6]. Hierarchical MARL further separated system operators and EV users, and coordinated them to minimize demand charges and energy costs [20].
Although literature in Sec. 5.1 and  5.2 optimized charging control with various aspects, most studies do not address the issue of uncertain EV departures. While some works[15, 8, 27] considered uncertainties in EV behavior, they primarily focused on a single objective, such as maximizing system revenue or minimizing user charging costs. In contrast, our work tackles dynamic factors, including fluctuating electricity prices and uncertain EV departures, aiming to simultaneously minimize total electricity costs for the charging system while accounting for EV users’ charging demands.
6 Conclusion
In this paper, we propose HUCA, a novel hierarchical multi-agent framework designed for real-time charging control in the workplace charging scenario. To address practical dynamics and limitations, HUCA integrates a hierarchical control and uncertainty-aware critic augmentation. With this design, HUCA adapts to dynamic factors, optimizes charging or discharging power for each charging pile, and accounts for departure uncertainties in evaluating the quality of charging control. Experiments demonstrate that, in both certain and uncertain departure scenarios, HUCA outperforms existing methods by achieving the lowest electricity costs while maintaining competitive performance in meeting EV charging requirements. A case study further illustrates that HUCA effectively balances energy supply between the building and EVs using real-time information.
A Reproducibility
A.1 Detailed Experimental Settings.
A.1.1 Dataset Preprocessing.
We utilize the CU-BEMS dataset [17], which contains detailed energy consumption data from an office building, recorded at one-minute intervals from July 2018 to December 2019. For dynamic pricing (real-time pricing) data, we use the APIs [5] provided by ComEd (Commonwealth Edison), an electric utility company, offering real-time 5-minute electricity pricing data through its hourly pricing program. In our study, we aggregate the total energy consumption from all electrical appliances in the CU-BEMS dataset into hourly data to represent the building’s energy consumption in our workplace charging scenario. The dynamic pricing data obtained from ComEd APIs is similarly aggregated into hourly intervals. The training period covers July 1, 2018, to August 31, 2018, and the testing period runs from September 1, 2018, to September 30, 2018.
A.1.2 Penalty Mechanism.
For the design of the penalty mechanism, the penalty charge is typically two or three times the basic capacity rate [18]. We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption)[7, 18], with detailed settings shown in Table 4. Let the maximum instant electricity load during the testing period be $L_{max}$. If $L_{max}<C_{contract}$, then the penalty is set to zero. Given a threshold coefficient $10<\delta<$ 1, the penalty is calculated by dividing the excess electricity into two tiers:
  1. 1.
@@ A Reproducibility
Based on this setting, the total penalty cost is estimated as follows:
(A.1) Penalty $\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
A.1.3 Hyperparamter Settings.

### v5

@@ 4 The HUCA Framework
However, early departures of EVs occur occasionally, making the departure time uncertain and complicating the selection of the best low-level actions. To tackle this challenge, inspired by the Upper Confidence Bound (UCB) algorithm [auer2002finite], which adjusts action-value estimations based on uncertainty (or confidence) and thereby affects the action decision, we propose a novel augmented action-value function, $\mathcal{Q}^{i}_{l}$, for the actor network of each low-level agent $c_{i}$.
[…]
Our uncertainty-aware critic augmentation is illustrated in Figure 3, and the augmented action-value function $\mathcal{Q}^{i}_{l}$ is designed as follows:
(4.6) $\hskip-8.5359pt\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})\cdot\biggl(1$-
@@
$\hskip 51.21504pt\rho\cdot\underbrace{|\log_{2}(a^{i}_{l}+\epsilon)|\cdot\sqrt{\frac{\max(\Delta SoC^{i}_{t},0)}{\Delta T^{i}_{t}}}}_{\text{the uncertainty factor}}\>\>\biggr)\Large\small\large,$
@@
where $a^{i}_{l}\in[0,1]$ is the action chosen by low-level agent $c_{i}$. A smaller $a^{i}$ indicates a tendency for higher discharge power. $\Delta SoC^{i}_{t}$ represents the difference between the expected SoC $SoC^{i}_{\text{dep}}$ and the current SoC $SoC^{i}_{t}$ of the EV connected to low-level agent (pile) $c_{i}$. $\Delta T^{i}_{t}$ is the time difference between the current time $t$ and the planned departure time $t^{i}_{\text{dep}}$. $\rho\in\mathbb{R}$ is a fixed coefficient representing the impact of uncertainty, and $\epsilon$ is a small constant. Consequently, high power discharging action is penalized when the uncertainty factor increases.
@@
Based on Eq. 4.6, the actor network of low-level agent $c_{i}$, with parameters $\theta^{\mu,i}_{l}$, updates the low-level policy $\mu^{i}_{l}$ with gradient descent as:
@@
(4.7) $\small\nabla_{\theta^{\mu,i}_{l}}J(\mu^{i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a}\sim\mathcal{D}_{l}}\bigl[\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(a^{i}_{l}|s^{i}_{l})\nabla_{a^{i}_{h}}\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})|_{a^{i}_{l}=\mu^{i}_{l}(s^{i}_{l})}\bigr].$}
Similar to the high-level agent, each low-level agent is trained with mini-batch updates. At every training step, a low-level agent $c_{i}$ draws a mini-batch $\mathcal{B}_{l}$ of transitions $\bigl\{(\textbf{x}^{(b)},a^{1,(b)}_{l},...,a^{N,(b)}_{l},r^{1,(b)}_{l},...,r^{N,(b)}_{l},\textbf{x}^{\prime(b)})\bigr\}_{b=1}^{\mathcal{B}_{l}}$ from the low-level replay buffer $\mathcal{D}_{l}$. The corresponding critic and actor networks, parameterized by $\theta^{Q,i}_{l}$ and $\theta^{\mu,i}_{l}$, are then updated as described in Algorithm 2.
Algorithm 2 Low-Level Agent Update
1:procedure UpdateLow($\mathcal{B}_{l}$, $\theta^{Q,i}_{l}$, $\theta^{\mu,i}_{l}$)
2:  Update the critic network by minimizing: $\mathcal{L}(\theta^{Q,i}_{l})=\frac{1}{\mathcal{B}_{l}}\sum_{b}\Big(Q^{i}_{l}(\textbf{x}^{(b)},a^{1,(b)}_{l},...,a^{N,(b)}_{l})-y^{i,(b)}_{l}\Big)^{2}$
3:
4:  Compute the augmented action-value function $\mathcal{Q}^{i}_{l}$ $\triangleright$ Eq. 4.6
5:  Update the actor network based on $\mathcal{Q}^{i}_{l}$: $\nabla_{\theta^{\mu,i}_{l}}J\approx\frac{1}{\mathcal{B}_{l}}\sum_{b}\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(s^{i,(b)}_{l})\nabla_{a^{i,(b)}_{l}}\mathcal{Q}^{i}_{l}(\textbf{x}^{(b)},\textbf{a}^{(b)})$
6:end procedure
@@
Continuous actions and rewards for low-level agents. In Eq. 4.6, the action $a^{i}_{l}$ is derived from the raw logits $g^{i}_{l}$ produced by the low-level policy $\mu^{i}_{l}$ of agent $c_{i}$. To further entangle low-level action space with the high-level discrete decision $a^{\text{disc}}_{h}$, the continuous action $a^{i}_{l}$ is formulated as follows:
@@
(4.8) $a^{i}_{l}=\left\{\begin{matrix}\delta+\delta\cdot\sigma(g^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=1$}\\ \delta\cdot\sigma(g^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=0$}\end{matrix}\right$.,
@@
(4.9) $\mathcal{P}^{\text{opt}}_{i,t}=a^{i}_{l}\cdot(\mathcal{P}^{\max}_{i,t}-\mathcal{P}^{\min}_{i,t})+\mathcal{P}^{\min}_{i,t}.$
Based on the optimal power decision $\mathcal{P}^{\text{opt}}_{i,t}$ at each time slot, the reward $r^{i}_{l}$ for the low-level agent (pile) $c_{i}$ is formulated considering the charging cost at the current time slot $t$ and the difference between the current SoC and the expected SoC at the EV’s departure connected to $c_{i}$, as defined below:
(4.10) $r^{i}_{l}=\omega\cdot(-\mathcal{P}^{\text{opt}}_{i,t}\cdot p_{t})+(-|SoC^{i}_{t}-SoC^{i}_{\text{dep}}|),$
where $p_{t}$ and $SoC^{i}_{t}$ are the energy price and the SoC of the EV at time slot $t$, respectively. $SoC^{i}_{\text{dep}}$ is the expected SoC at the EV’s departure. The parameter $\omega$ reflects the impact of the current charging cost. This reward design encourages low-level agents to discharge EVs when the energy price is high, while also aiming to meet their expected SoC.
@@
In Secs. 4.1 and 4.2, the critic and actor networks are optimized using mini-batches sampled from the replay buffers. To stabilize training, a soft target-network update is applied. Let $\theta_{h}=\{\theta^{Q}_{h},\theta^{\mu}_{h}\}$ and $\theta^{i}_{l}=\{\theta^{Q,i}_{l},\theta^{\mu,i}_{l}\}$ denote the parameters of the high-level and the $i$-th low-level networks, respectively, and let $\theta_{h}^{\prime}=\{\theta^{\prime Q}_{h},\theta^{\prime\mu}_{h}\}$ and $\theta^{{}^{\prime}i}_{l}=\{\theta^{\prime Q,i}_{l},\theta^{\prime\mu,i}_{l}\}$ be their corresponding target network parameters. These target parameters are updated as follows:
@@
where $\tau\in[0,1]$ is the coefficient controlling the update rate.
The overall architecture of HUCA is shown in Algorithm 3. This design enables both high-level and low-level agents to learn optimal charging control policies by considering dynamic factors. Therefore, HUCA provides more effective real-time charging control without requiring future information.
Algorithm 3 The HUCA framework
1:The maximum learning episode $E$; charging piles $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$; EV set $\mathcal{V}_{t}$; EV charging information $\{\mathcal{I}^{i}|v_{i}\in\mathcal{V}_{t}\}$; power boundaries $\{\mathcal{P}^{\text{max}}_{i,t},\mathcal{P}^{\text{min}}_{i,t}|v_{i}\in\mathcal{V}_{t}\}$.
2:Optimal policies $\mu_{h}$ and $\{\mu_{l}^{\,i}\}_{i=1}^{N}$ with parameters $\theta^{\mu}_{h}$ and $\{\theta^{\mu,i}_{l}\}_{i=1}^{N}$
3:Initialize parameters $\theta^{\mu}_{h},\theta^{Q}_{h}$ for the high-level agent and $\{\theta^{\mu,i}_{l},\theta^{Q,i}_{l}\}_{i=1}^{N}$ for the low-level agents; replay buffers $\mathcal{D}_{h},\mathcal{D}_{l}$.
4:for episode = 1 to $E$ do
5:  Receive initial state $s_{h}$, $\textbf{x}=\{s^{i}_{l}\}_{i=1}^{N}$
6:  for $t=1$ to max-episode-length do
7:    // Select actions
8:    Select and compute action $a^{disc}_{h}$
9:    for each low-level agent i=1,,Ni=1,...,N do
10:     Select and compute action $a^{i}_{l}$ based on $a^{disc}_{h}$
11:     Estimate optimal power $\mathcal{P}^{\text{opt}}_{i,t}$
12:    end for
13:    // Store transition experiences
14:    Apply optimal powers $\{\mathcal{P}^{\text{opt}}_{i,t}\}_{i=1}^{N}$ to EVs and receive rewards $(r_{h},\{r^{i}_{l}\}_{i=1}^{N})$
15:    Observe next states $(s^{\prime}_{h},\textbf{x}^{\prime}=\{s^{\prime i}_{l}\}_{i=1}^{N})$
16:    Store $(s_{h},a_{h},r_{h},s^{\prime}_{h})$ in $\mathcal{D}_{h}$
17:    Store $(\textbf{x},a^{1}_{l},...,a^{N}_{l},r^{1}_{l},...,r^{N}_{l},\textbf{x}^{\prime})$ in $\mathcal{D}_{l}$
18:    Update current states
19:  end for
20:  // Update networks
21:  Sample a mini-batch $\mathcal{B}_{h}$ from $\mathcal{\mathcal{}}{D}_{h}$
22:  UpdateHigh($\mathcal{B}_{h}$, $\theta^{Q}_{h}$, $\theta^{\mu}_{h}$) $\triangleright$ Algorithm 1
23:  Sample a mini-batch $\mathcal{B}_{l}$ from $\mathcal{D}_{l}$
24:  for each low-level agent i=1,,Ni=1,...,N do
25:    UpdateLow($\mathcal{B}_{l}$, $\theta^{Q,i}_{l}$, $\theta^{\mu,i}_{l}$) $\triangleright$ Algorithm 2
26:  end for
27:  Soft target networks update
28:end for
@@ 5 Experimental Results
5.1 Dataset and Experimental Setup
5.1.1 Dataset and Preprocessing.
@@ 5.1 Dataset and Experimental Setup
For our workplace charging scenario, following datasets are used: (i) Building electricity demands: the CU-BEMS dataset [Pipattanasomporn2020] provides office building electricity consumption recorded at one-minute intervals from July 2018 to December 2019. (ii) Pricing mechanism: we use the ComEd APIs [ComEdHourlyPricing], which provide real-time 5-minute electricity pricing data through its hourly pricing program. All data are aggregated into hourly intervals. The training period spans July 1 to August 31, 2018, with testing conducted in the following month.
5.1.2 Penalty Mechanism.
The penalty charge is typically two or three times the basic capacity rate [rosado2020framework]. We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption) [Fernndez2013CostOO] [rosado2020framework], with detailed settings shown in Table 1. Let the maximum instant electricity load during the testing period be $L_{max}$. If $L_{max}<C_{contract}$, then the penalty is set to zero. Given a threshold coefficient $0<\delta<1$, the penalty is calculated by dividing the excess electricity into two tiers:
  1. (a)
@@
Based on this setting, the total penalty cost is estimated as:
(5.12) $\hskip 0.0pt\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
@@
Table 2: Random variables for EV information.
Information Distribution Boundaries
Arrival time $\mathcal{N}(9,1^{2})$ $7\leq t^{i}_{\text{arr}}\leq 12$
[…]
Certain Departure Scenario
Metrics CP Num. Methods
OPT DDPG IQL VDN MADDPG HUCA (ours)
Penalty Cost ($\downarrow$) 10 0.00(▲100.0%) 3682.85(▽937.9%) 2516.84(▽609.3%) 522.28(▽47.2%) 354.84(▽0.0%) 354.82
Total Cost ($\downarrow$) 5805.28(▲6.6%) 9563.31(▽53.9%) 8426.36(▽35.6%) 6431.31(▽3.5%) 6220.33(▽0.1%) 6215.42
Penalty Cost ($\downarrow$) 20 0.00(▲100.0%) 7818.52(▽375.7%) 5643.15(▽243.4%) 2271.32(▽38.2%) 1869.38(▽13.7%) 1643.47
Total Cost ($\downarrow$) 5860.05(▲24.2%) 13929.11(▽80.2%) 11798.68(▽52.7%) 8413.61(▽8.9%) 7952.69(▽2.9%) 7728.57
Uncertain Departure Scenario
Metrics CP Num. Methods
OPT DDPG IQL VDN MADDPG HUCA (ours)
Penalty Cost ($\downarrow$) 10 0.00(▲100.0%) 4671.74(▽1212.9%) 1742.35(▽389.6%) 529.18(▽48.7%) 448.52(▽26.0%) 355.84
Total Cost ($\downarrow$) 5691.96(▲5.0%) 10398.77(▽73.5%) 7479.04(▽24.8%) 6214.53(▽3.7%) 6087.48(▽1.6%) 5993.78
[…]
Penalty Cost ($\downarrow$) 20 0.00(▲100.0%) 5010.81(▽1311.6%) 2002.12(▽464.1%) 966.51(▽172.3%) 931.01(▽162.3%) 354.90
Total Cost ($\downarrow$) 5760.05(▲4.2%) 10823.07(▽80.0%) 7815.99(▽30.0%) 6721.82(▽11.8%) 6562.80(▽9.1%) 6013.52
[…]
5.1.3 Simulation of EV Behaviors.
The charging information $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$ (Def. 2) of EV $v_{i}$ is modeled using normal distributions by following [LiConstrainedLR]. Table 2 shows the setting of random variables and $C^{i}$ is set to 60 kWh. Two departure scenarios for determining the actual departure time $\widehat{t}^{i}_{dep}$ of EV $v_{i}$ are examined:
  1. (a)
    Certain departure scenario: $\widehat{t}^{i}_{dep}$ is the same as the expected departure time $t^{i}_{dep}$.
  2. (b)
    Uncertain departure scenario: $\widehat{t}^{i}_{dep}$ is randomly sampled earlier than the expected departure time, $\widehat{t}^{i}_{dep}\in[1,t^{i}_{dep})$.
5.1.4 Baselines.
We compare our HUCA against several baselines, including an oracle with full knowledge of future information, the single-agent model(DDPG [lillicrap2015continuous]) and multi-agent reinforcement learning models (IQL [Tan1997MultiAgentRL], VDN[Sunehag2017ValueDecompositionNF], and MADDPG [lowe2017multi]), as described below:
  • DDPG (Deep Deterministic Policy Gradient) [lillicrap2015continuous]: An actor-critic algorithm designed for continuous control in single-agent environments.
  • IQL (Independent Q-Learning) [Tan1997MultiAgentRL]: A decentralized multi-agent reinforcement learning approach where each agent independently learns its own Q-function.
  • VDN (Value Decomposition Networks) [Sunehag2017ValueDecompositionNF]: A centralized value-based reinforcement learning method used in multi-agent systems.
  • MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [lowe2017multi]: Extends DDPG to multi-agent settings by incorporating centralized training and decentralized execution.
5.1.5 Evaluation Metrics.
Following metrics are used to evaluate performance:
  • Penalty Cost (USD) measures the cost of exceeding the contracted capacity [Fernndez2013CostOO], which multiplies the penalty rate with the peak amount of electricity. In accordance with [rosado2020framework], the penalty mechanism is designed as detailed in Table 1.
5.1.6 Hyperparamter Settings.